# Projected Density Matrix Sampling for Lattice Hamiltonians

### v1

@@ VI.6 The Quantum Heisenberg Limit
We now investigate the performance of our pdms method on an $L=8$ lattice, which lies beyond the reach of exact diagonalization studies. We first focus on the case $(J=0,J^{\prime}=1)$, which turns the SSM into the well studied two dimensional quantum Heisenberg model. In this study we focus on the $\mathbb{P}_{+}$ subspace. In the quantum Heisenberg limit, we can show that $Z_{11}/Z^{b}=1$ for all values of $\beta$ and $L$, a result that can be easily proved, for example, using meron-cluster properties [20]. This implies that the sign problem is absent, at least for the projection state $|\psi_{1}\rangle$ at this coupling.
@@
In this limit, our Monte Carlo method therefore enables a high-precision determination of $e_{1}(\beta)$ associated with the projection subspace $\mathbb{P}_{(1)}$ spanned by $|\psi_{1}\rangle$. At large values of $\beta$, this projected thermal energy is expected to yield the ground-state energy of the two-dimensional Heisenberg antiferromagnet. Results for the $L=8$ lattice are shown in the second column of Table 6 and also plotted in Fig. 9. We observe that $e_{1}(\beta)$ decreases gradually with increasing $\beta$. A previous study performed at $\beta=64$ reported the ground-state energy to be 43.103168(250)-43.103168(250) [35], whereas our current calculation yields, $e_{1}(\beta=15)=-43.0724(3)$, which has clearly not converged to the low temperature limit.
@@
Rather than taking the large-$\beta$ limit with $\mathbb{P}_{(1)}$, we can instead use $\mathbb{P}_{+}$ to test whether increasing the dimension of the projection subspace improves the rate of convergence. The projected thermal energies obtained using $\mathbb{P}_{+}$ are listed in the right four columns of Table 6 and also plotted in Fig. 9. We now find $e_{1}(\beta=3.0)=-43.11(6)$ and $e_{1}(\beta=5.0)=-43.11(1)$, both showing clear signs of convergence to the ground state result of [35]. The fact that we can obtain the ground state energy with an error of less than 0.20.2% even at the value of $\beta=3.0$ is impressive.
@@
We note that the Monte Carlo errors in the calculation of $e_{1}(\beta)$ arise from contributions across the entire projection subspace. Sign problems can reemerge through other projection states for which $Z_{aa}/Z^{b}$ becomes small as $\beta$ increases. The use of larger projection subspaces therefore typically restricts us to smaller values of $\beta$. We show this behavior in Table 6. At $\beta=3.0$, we are able to extract all four $e_{a}(\beta)$ values, but at $\beta=5.0$ we had to discard the state $|\psi_{4}\rangle$ to obtain reliable results due to zero modes creeping in from that state. However, this reduction still allows us to recover the three lowest levels with equal or higher accuracy. Notably, ${d_{\mathrm{proj}}}=3$ performs better than ${d_{\mathrm{proj}}}=1$ for the ground-state calculation at $\beta=5.0$. For $\beta=10$ and $15$, the projection subspace had to be reduced further. By retaining only $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, we obtained more accurate estimates for the ground-state energy, 43.1036(3)-43.1036(3), and the first excited-state energy, 42.927(1)-42.927(1). This estimate of the ground-state energy is consistent with that obtained at $\beta=5.0$ with ${d_{\mathrm{proj}}}=3$. Another earlier work (see [36]) reported the first excited energy as 42.90(4)-42.90(4), consistent with our result but with substantially larger errors. For the second excited state, the previous estimate of 42.54(3)-42.54(3) and our result of 42.63(8)-42.63(8) at $\beta=3$ are also consistent, although our uncertainties are larger. We expect that a more carefully chosen projection subspace could improve the precision of some the higher excited states further.
@@ VI.7 Perturbative Energy Gap at small $J^{\prime}/J$
In Section VI.2 we explained that $\mathbb{P}_{-}$ is an optimal subspace for computing the first excited energy level for small values of $J^{\prime}/J$. Here we verify this by exploring how the $\mathbb{P}_{-}$ subspace performs at $(J=1,J^{\prime}=0.5)$. The energy gap between the ground state ($E_{1}$) and the first excited state ($E_{2}$) was computed in perturbation theory to very high orders [37] and the first few terms of this series are given by
@@ VII Conclusions
Using these analytic insights, we demonstrated in Sections V and VI that our method remains robust across a wide range of projection subspaces at high temperatures. In practice, we find that stable convergence close to the projected low-energy spectrum can often be achieved at moderate values of $\beta$, without encountering severe sign problems. At larger $\beta$ or with significantly expanded projection subspaces, sign problems may reemerge, similar to the challenges faced when extracting higher-energy excitations in otherwise sign-problem-free systems [38]. These observations highlight the importance of carefully selecting projection subspaces that target the desired low-energy levels efficiently. We demonstrated this strategy successfully in our study of the transverse-field Ising model and $J=0$ regime of the SSM, where well-chosen projection states enabled rapid and stable convergence.
A further strength of the pdms framework is that the Monte Carlo sampling can, in principle, be formulated in a variety of bases within the Hilbert space. In some cases, sign problems can be completely resolved through a careful choice of basis [34, 39], while in others they can be significantly mitigated [40]. It may also be possible to take guidance from analytic techniques such as perturbation theory when constructing the sampling basis. Furthermore, new ideas for addressing sign problems, such as meron clusters [20], fermion bags [41], exponential error reduction techniques using the multi-level blocking ideas [42, 43, 44], can all be naturally incorporated into our framework. These features showcase the broad adaptability of the method and its potential for integration with parallel advances in Monte Carlo methodology.
An interesting avenue for further development is the use of physically-motivated construction of projection states that are guided by the known quantum phases of the underlying Hamiltonian, as demonstrated in recent tensor-network studies [23]. For example, projection subspaces inspired by phase-specific wavefunctions, such as plaquette-valence-bond (PVB) type states in the Shastry–Sutherland model, may provide more efficient starting points for our pdms method. Our preliminary results show that such physically motivated choices may considerably accelerate convergence and improve stability in challenging regimes. We plan to share these findings in a future work.
@@ Appendix A Stochastic Series Expansion
While we chose to present our pdms algorithm using a path integral formulation, it can equally well be formulated using the popular stochastic series expansion (SSE) [16]. The steps are self evident but are included here for completeness. To compute the $Z$ matrix, we simply expand the exponential as usual in the SSE,

### v2

@@ VI.6 The Quantum Heisenberg Limit
We now investigate the performance of our pdms method on an $L=8$ lattice, which lies beyond the reach of exact diagonalization studies. We first focus on the case $(J=0,J^{\prime}=1)$, which turns the SSM into the well studied two dimensional quantum Heisenberg model. In this study we focus on the $\mathbb{P}_{+}$ subspace. In the quantum Heisenberg limit, we can show that $Z_{11}/Z^{b}=1$ for all values of $\beta$ and $L$, a result that can be easily proved, for example, using meron-cluster properties [Chandrasekharan:1999cm]. This implies that the sign problem is absent, at least for the projection state $|\psi_{1}\rangle$ at this coupling.
@@
In this limit, our Monte Carlo method therefore enables a high-precision determination of $e_{1}(\beta)$ associated with the projection subspace $\mathbb{P}_{(1)}$ spanned by $|\psi_{1}\rangle$. At large values of $\beta$, this projected thermal energy is expected to yield the ground-state energy of the two-dimensional Heisenberg antiferromagnet. Results for the $L=8$ lattice are shown in the second column of Table 6 and also plotted in Fig. 9. We observe that $e_{1}(\beta)$ decreases gradually with increasing $\beta$. A previous study performed at $\beta=64$ reported the ground-state energy to be 43.103168(250)-43.103168(250) [PhysRevB.56.11678], whereas our current calculation yields, $e_{1}(\beta=15)=-43.0724(3)$, which has clearly not converged to the low temperature limit.
@@
Rather than taking the large-$\beta$ limit with $\mathbb{P}_{(1)}$, we can instead use $\mathbb{P}_{+}$ to test whether increasing the dimension of the projection subspace improves the rate of convergence. The projected thermal energies obtained using $\mathbb{P}_{+}$ are listed in the right four columns of Table 6 and also plotted in Fig. 9. We now find $e_{1}(\beta=3.0)=-43.11(6)$ and $e_{1}(\beta=5.0)=-43.11(1)$, both showing clear signs of convergence to the ground state result of [PhysRevB.56.11678]. The fact that we can obtain the ground state energy with an error of less than 0.20.2% even at the value of $\beta=3.0$ is impressive.
@@
We note that the Monte Carlo errors in the calculation of $e_{1}(\beta)$ arise from contributions across the entire projection subspace. Sign problems can reemerge through other projection states for which $Z_{aa}/Z^{b}$ becomes small as $\beta$ increases. The use of larger projection subspaces therefore typically restricts us to smaller values of $\beta$. We show this behavior in Table 6. At $\beta=3.0$, we are able to extract all four $e_{a}(\beta)$ values, but at $\beta=5.0$ we had to discard the state $|\psi_{4}\rangle$ to obtain reliable results due to zero modes creeping in from that state. However, this reduction still allows us to recover the three lowest levels with equal or higher accuracy. Notably, ${d_{\mathrm{proj}}}=3$ performs better than ${d_{\mathrm{proj}}}=1$ for the ground-state calculation at $\beta=5.0$. For $\beta=10$ and $15$, the projection subspace had to be reduced further. By retaining only $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, we obtained more accurate estimates for the ground-state energy, 43.1036(3)-43.1036(3), and the first excited-state energy, 42.927(1)-42.927(1). This estimate of the ground-state energy is consistent with that obtained at $\beta=5.0$ with ${d_{\mathrm{proj}}}=3$. Another earlier work (see [PhysRevB.49.5983]) reported the first excited energy as 42.90(4)-42.90(4), consistent with our result but with substantially larger errors. For the second excited state, the previous estimate of 42.54(3)-42.54(3) and our result of 42.63(8)-42.63(8) at $\beta=3$ are also consistent, although our uncertainties are larger. We expect that a more carefully chosen projection subspace could improve the precision of some the higher excited states further.
@@ VI.7 Perturbative Energy Gap at small $J^{\prime}/J$
In Section VI.2 we explained that $\mathbb{P}_{-}$ is an optimal subspace for computing the first excited energy level for small values of $J^{\prime}/J$. Here we verify this by exploring how the $\mathbb{P}_{-}$ subspace performs at $(J=1,J^{\prime}=0.5)$. The energy gap between the ground state ($E_{1}$) and the first excited state ($E_{2}$) was computed in perturbation theory to very high orders [PhysRevB.65.014408] and the first few terms of this series are given by
@@ VII Conclusions
Using these analytic insights, we demonstrated in Sections V and VI that our method remains robust across a wide range of projection subspaces at high temperatures. In practice, we find that stable convergence close to the projected low-energy spectrum can often be achieved at moderate values of $\beta$, without encountering severe sign problems. At larger $\beta$ or with significantly expanded projection subspaces, sign problems may reemerge, similar to the challenges faced when extracting higher-energy excitations in otherwise sign-problem-free systems [Wagman:2017xfh]. These observations highlight the importance of carefully selecting projection subspaces that target the desired low-energy levels efficiently. We demonstrated this strategy successfully in our study of the transverse-field Ising model and $J=0$ regime of the SSM, where well-chosen projection states enabled rapid and stable convergence.
A further strength of the pdms framework is that the Monte Carlo sampling can, in principle, be formulated in a variety of bases within the Hilbert space. In some cases, sign problems can be completely resolved through a careful choice of basis [PhysRevB.57.R3197, PhysRevLett.117.197203], while in others they can be significantly mitigated [JensEisert2020]. It may also be possible to take guidance from analytic techniques such as perturbation theory when constructing the sampling basis. Furthermore, new ideas for addressing sign problems, such as meron clusters [Chandrasekharan:1999cm], fermion bags [Chandrasekharan:2013rpa], exponential error reduction techniques using the multi-level blocking ideas [Luscher:2001up, Mak:1998zw, Yoo:2004mh], can all be naturally incorporated into our framework. These features showcase the broad adaptability of the method and its potential for integration with parallel advances in Monte Carlo methodology.
An interesting avenue for further development is the use of physically-motivated construction of projection states that are guided by the known quantum phases of the underlying Hamiltonian, as demonstrated in recent tensor-network studies [Corboz2025]. For example, projection subspaces inspired by phase-specific wavefunctions, such as plaquette-valence-bond (PVB) type states in the Shastry–Sutherland model, may provide more efficient starting points for our pdms method. Our preliminary results show that such physically motivated choices may considerably accelerate convergence and improve stability in challenging regimes. We plan to share these findings in a future work.
@@ Appendix A Stochastic Series Expansion
While we chose to present our pdms algorithm using a path integral formulation, it can equally well be formulated using the popular stochastic series expansion (SSE) [Sandvik2010_QMCReview]. The steps are self evident but are included here for completeness. To compute the $Z$ matrix, we simply expand the exponential as usual in the SSE,